OpenStudy (anonymous):

Suppose the integral from 2 to 6 of g of x, dx equals 12 and the integral from 5 to 6 of g of x, dx equals negative 3, find the value of the integral from 2 to 5 of 3 times g of x, dx .

OpenStudy (anonymous):

https://gyazo.com/d93ca0e3dc7d22cf4cb98fb0d12da039

OpenStudy (anonymous):

@jim_thompson5910 Would it not be -9?

OpenStudy (caozeyuan):

Should be -9, I am really confused

OpenStudy (anonymous):

Same Unless there's another method

OpenStudy (tkhunny):

I can't read your link. Recommendation: Learn a little LaTeX and give it a good show. Generally, \(\int\limits^{6}_{2}g(x)\;dx = \int\limits^{5}_{2}g(x)\;dx + \int\limits^{6}_{5}g(x)\;dx\)

OpenStudy (anonymous):

@tkhunny that is not what they ask for. They have \[\int\limits_{2}^{6}g(x)= 12 and \int\limits_{5}^{6}g(x)=-3 what is \int\limits_{5}^{6}3g(x)\]

jimthompson5910 (jim_thompson5910):

@Ephemera you wrote `find the value of the integral from 2 to 5 of 3 times g of x, dx .` for the third integral but the third integral shown in the link is \[\Large \int_{5}^{6}3g(x)dx\]

OpenStudy (anonymous):

That's weird I copied and pasted the question and it seems the pictures translated into text. So there must be an error in the photo.

OpenStudy (anonymous):

Thanks for pointing that out.

jimthompson5910 (jim_thompson5910):

yeah I'm guessing they meant to write \[\Large \int_{2}^{5}3g(x)dx\]

OpenStudy (anonymous):

I got 45

jimthompson5910 (jim_thompson5910):

use the equation @tkhunny wrote out and you can isolate \[\Large \int_{2}^{5}3g(x)dx\]think of it as a variable

OpenStudy (anonymous):

That's what was done. Got 15 for 2 to 5 then multiplied by 3 to get 45

jimthompson5910 (jim_thompson5910):

oh sorry, without the 3, but yeah the final answer will be 45

OpenStudy (anonymous):

Mind helping with another? This one really has me puzzled.

jimthompson5910 (jim_thompson5910):

sure